Part A Constants | Periodic Table The California sea lion is capable of making extremely fast, tight turns while swimming underwater. In one study, scientists observed a sea lion making a circular turn with a radius of 0.31 m while swimming at 4.4 m/s. What is the sea lions centripetal acceleration, in units of g? Express your answer in units of 9. ×g Submit Request Answer PartB What percentage is this acceleration of that of an F-15 fighter jets maximum centripetal acceleration of 9g? Express your answer as a percentage
Answer #1

31 4.4 m 9-81 2 - (4.4 o 31 6 2 -4512 ac = 62-451 9 8 = (6-366 ) 6-366 70-73 %